# algorithmic

### v1

@@ 9 Implication for Practical Applications
For link prediction on dense graphs, the contribution of node embeddings becomes dominant. Simple GNNs like GCN[29] with the incorporation of trainable node embeddings can outperform most existing methods, rendering such a solution an optimal choice. However, this does not mean that the model using node embeddings will certainly perform better than that only using node features, especially in practical applications where careful feature engineering guided by domain knowledge is conducted. Besides, the use of trainable node embeddings remains limitations in the inductive setting [60], where new nodes are added to the graph, and the model together with all node embeddings may need to be retrained. In such cases, the methods that do not involve the training of node embeddings may offer more practical suitability.
10 Limitations
@@
11 Conclusion
Link prediction stands as a pivotal task within the realm of graph applications. Our exploration into this domain reveals noteworthy variations in the performance of various link prediction methods across different graphs, with a significant dependence on graph degrees. Notably, on dense graphs, we observe that straightforward GNNs, like GCN, exhibit superior link prediction performance compared to many models that are developed specifically for link prediction. In contrast, on sparse graphs, the simple common-neighbor method often outshines GNN-based approaches. Understanding and interpreting these performance fluctuations is imperative, serving as a compass for refining existing methodologies and establishing a foundation for the development of more effective link prediction algorithms.

### v2

@@ IX Implication for Practical Applications
For link prediction on dense graphs, the contribution of node embeddings becomes dominant. Simple GNNs like GCN[kipf2017semi] with the incorporation of trainable node embeddings can outperform most existing methods, rendering such a solution an optimal choice. However, this does not mean that the model using node embeddings will certainly perform better than that only using node features, especially in practical applications where careful feature engineering guided by domain knowledge is conducted. Besides, the use of trainable node embeddings remains limitations in the inductive setting [teru2020inductive], where new nodes are added to the graph, and the model together with all node embeddings may need to be retrained. In such cases, the methods that do not involve the training of node embeddings may offer more practical suitability.
X Limitations
@@
XI Conclusion
Link prediction stands as a pivotal task within the realm of graph applications. Our exploration into this domain reveals noteworthy variations in the performance of various link prediction methods across different graphs, with a significant dependence on graph degrees. Notably, on dense graphs, we observe that straightforward GNNs, like GCN, exhibit superior link prediction performance compared to many models that are developed specifically for link prediction. In contrast, on sparse graphs, the simple common-neighbor method often outshines GNN-based approaches. Understanding and interpreting these performance fluctuations, which is the objective of this work, is imperative, serving as a compass for refining existing methodologies and establishing a foundation for the development of more effective link prediction algorithms.